OpenStudy (anonymous):

find all possible rational zeros of the polynomial function f(x) = 4x^3 – 5x^2 + 9x – 6. I did it and I got +-1 +-2+- 3+- 6 +-1/2 +- 3/2 +- 1/4+- 2/4+- 3/4+- 6/4 can some one please check my work thanks

OpenStudy (anonymous):

hold up

OpenStudy (anonymous):

Try pluging all values till 6 and -6, it doenst look like this function have rational zeros.

OpenStudy (anonymous):

All possible rational zeros: p/q where p is the constant and q is the leading coefficient. ±1/4, ±1/2, ±3/4, ±1, ±3/2, ±2, ±3, ±6.

OpenStudy (anonymous):

±2/4 and ±6/4 are already there, but simplified.

OpenStudy (anonymous):

can you explain to me what your tallking about please

OpenStudy (anonymous):

Alright, Do you know the rational root theorem though? How you got the ±1, ±2, etc?

OpenStudy (anonymous):

First I found the factors of -6 and 4

OpenStudy (anonymous):

Yeah. THat's called the rational root theorem. Your answer is generally correct, but all I was saying was that 6/4 = 3/2 and 2/4 = 1/2 which are already there. Those additional ones like I said are unnecessary.

OpenStudy (anonymous):

So, your answer is correct :)

OpenStudy (anonymous):

okay just write out the final answer for me to make sure

OpenStudy (anonymous):

±1/4, ±1/2, ±3/4, ±1, ±3/2, ±2, ±3, ±6